the length of a rectangular flower garden is 6 feet more than three times the width. the perimeter of the garden is 32 feet. what is the area of the garden?

What I did: 6+3n=w or 2l+2w=32 ..... Than I don't know.

Answer by rothauserc(4718)

length (l) = 3*width(w) + 6
2*(3w+6) + 2w = 32
6w+12 +2w = 32
8w = 20
w = 20/8 = 5/2 = 2.5 feet
l = 3*(2.5) + 6 = 13.5 feet
Area (A) = (13.5) * (2.5) = 33.75 square feet